YoursTruly:

Solve using any method, show all work, and round your answer to 2 decimal places if needed. x^2+8x+16=0

JustSaiyan:

Factor the left side first: (x+4)(x+4)=0 Set the factors equal, or to zero: x+4=0 or x+4=0 Answer: x=-4